# Nonsmooth Calabi–Yau structures for algebras and coalgebras

### v1

@@ 5.3. Properness for coalgebras
Let $M\in\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)$ be an acyclic comodule. We wish to show that $M$ is coacyclic. Following [56, 5.5], we see that the full subcategory of $D^{\mathrm{co}}(C)$ consisting of the acyclic $C$-comodules is precisely the left orthogonal to $C$. But this is also the left orthogonal to $\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)$. So $M$ is simultaneously an object of both $\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)$ and its left orthogonal, and hence is zero in $D^{\mathrm{co}}(C)$, as required. ∎
@@
There is a natural localisation functor $D^{\mathrm{co}}(C)\to D({\mathbf{Comod}}\text{-}C)$ whose kernel consists of the acyclic comodules. Under the assumption of Vopěnka’s principle, there is a natural equivalence $\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)\simeq D({\mathbf{Comod}}\text{-}C)$ so that $\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)$ is actually the largest subcategory of $D^{\mathrm{co}}(C)$ for which the natural functor to $D(k)$ reflects isomorphisms. See [56, 5.5, Note added three years later] for further discussion.
@@ 6.2. Calabi–Yau algebras
The original definition of Ginzburg CY in [23] requires in addition that the isomorphism $f:A\to A^{\vee}[n]$ in the derived category of $A$-bimodules satisfies the extra condition that the composition $A\to A^{\vee\vee}\xrightarrow{f^{\vee}[n]}A^{\vee}[n]$ is equal to $f$. However, it is a theorem of Van den Bergh that in the smooth setting, this extra condition is superfluous [66, C.1]. In the nonsmooth setting, this extra condition does not seem useful.
@@ 6.3. Calabi–Yau conditions on categories
If $A$ is Ginzburg $n$-CY then the pair $(D(A),{{\mathbf{pvd}}}\kern 1.0pt(A))$ has a $n$-CY structure [41, 3.4]. If $A$ is smooth, proper, and finitely presented, then $[n]$ is a Serre functor on ${{\mathbf{pvd}}}\kern 1.0pt(A)$ if and only if $A$ is Ginzburg $n$CY [23, 3.2.4].
@@
We follow the proof of [41, 4.1]. There are natural maps
@@ 7.2. Discrete Gorenstein algebras
Let $R$ be a commutative noetherian $k$-algebra of finite Krull dimension $d$. Recall from [37, 38] that $R$ is said to be singular CY if it is Gorenstein and equicodimensional [37, 3.10]. An algebra essentially of finite type is noetherian and of finite Krull dimension, and an irreducible algebra is equicodimensional. Hence, if $R$ is irreducible and essentially of finite type, then it is singular CY if and only if it is Gorenstein, and 7.8 provides implications
@@ 7.3. Frobenius algebras
Recall that a discrete finite dimensional graded algebra $A$ is $d$-Frobenius if there is an isomorphism $A^{*}\cong A[d]$ of $A$-modules (equivalently, $A^{\circ}$-modules). Such an algebra is equipped with a Nakayama automorphism $\nu$, and one has an $A$-bimodule isomorphism $A^{*}\cong A_{\nu}[d]$, where $A_{\nu}$ denotes $A$ with the usual bimodule structure on the left, and twisted by $\nu$ on the right. One then obtains a quasi-isomorphism $\operatorname{HH}_{\bullet}(A,A_{\nu})^{*}\simeq\operatorname{HH}^{\bullet}(A)[d]$. A symmetric Frobenius algebra is one whose Nakayama automorphism is the identity, so if $A$ is symmetric one has the simpler description $\operatorname{HH}_{\bullet}(A)^{*}\simeq\operatorname{HH}^{\bullet}(A)[d]$. For a comprehensive reference on the classical theory of Frobenius algebras, see [47, §16].
@@ 8.1. The main theorem
For all statements, the proof consists of matching up appropriate modules, co/contramodules, and functors, across the relevant version of Koszul duality. We begin with (1). By 3.9, across bimodule Koszul duality the diagonal bimodule $A$ corresponds to the diagonal bicomodule $C$. Hence by 3.14, across bimodule Koszul duality the inverse dualising complex $A^{\vee}$ corresponds to $D(C)$. So one is invertible if and only if the other is, which proves (1). The proof of (2) is identical.
@@
For (4), 3.1(6) shows that across one-sided Koszul duality the left $C$-contramodule $k$ corresponds to the left $A$-module $A^{*}$. Hence 3.17 shows that the left $C$-contramodule ${\mathrm{\mathbb{R}}}\mathrm{Hom}^{C}(k,C^{*})$ corresponds to the left $A$-module $A^{**}$. Hence if $A$ is $d$-Frobenius, we have $A^{*}\simeq A^{**}[d]$ as left $A$-modules and hence $C$ is $d$-Gorenstein. In the other direction, if $C$ is $d$-Gorenstein then the left $A$-module $M\coloneqq A^{*}$ is isomorphic to $M^{*}[d]$, which implies that $M\in{{\mathbf{ref}}}\kern 1.0pt{A}$ and hence $A\in{{\mathbf{ref}}}\kern 1.0pt{A}$. Hence $A^{**}\simeq A$ and so A is $d$-Frobenius.
For (5), 3.10 shows that across two-sided Koszul duality the bimodule $A^{*}$ corresponds to the bicontramodule $C^{*}$. Hence 3.19 implies that the $A$-bimodule $A^{**}$ corresponds to the $C$-bicontramodule $C^{\vee}$. If $A$ is twisted symmetric, we see that there is an invertible bimodule $X$ such that $A^{*}\simeq{}_{X}(A^{**})$. Translating this across Koszul duality, we see that there is an invertible bicomodule $Y$ such that $C^{*}\simeq{}_{Y}(C^{\vee})$; in other words $C$ is twisted CY. Conversely if $C$ is twisted CY then with we obtain an invertible $A$-bimodule $X$ and an $A$-bilinear quasi-isomorphism $A^{*}\simeq{}_{X}(A^{**})$, and so putting $M\coloneqq A^{*}$ we have $A$-bilinear quasi-isomorphisms
@@
and hence $R\otimes S$ is $(r+s)$-Gorenstein. For (2), suppose that $C$ is regular $r$-Frobenius and $D$ is regular $s$-Frobenius. By [33, 5.1] we have a natural algebra quasi-isomorphism $\Omega(C\otimes D)\simeq\Omega(C)\otimes\Omega(D)$. Putting $R\coloneqq\Omega C$ and $S\coloneqq\Omega D$ we see by 8.3 that $R$ is $r$-Gorenstein and $S$ is $s$-Gorenstein. Moreover $R$ is proper since $C$ is regular. Hence by (1), $R\otimes S$ is $(r+s)$-Gorenstein, and hence $C\otimes D$ is $(r+s)$-Frobenius by 8.3 again. ∎
@@ 8.2. Proper Frobenius coalgebras
Since $C$ is proper, it is weakly equivalent to a finite dimensional comodule over itself, and hence the comodule $C^{*}$ is the usual linear dual of $C$. By dualising, it follows that $A$ is $(-d)$-Frobenius if and only if $C$ is quasi-isomorphic to $C^{*}[d]$ as $C$-comodules. Since weak equivalences are quasi-isomorphisms, $(1)$ is now clear. The second claim follows similarly from 8.11. The third claims are proved identically; for (4) only needs to observe in addition that if $C$ is simply connected then so is $C^{e}$. ∎
@@
The simply connectedness assumptions of the above proposition cannot be dropped. Counterexamples can be found in chain coalgebras of topological spaces; see 9.41 for the details.
@@ 9.3. Hopf algebras
To begin the proof of 9.8, we need to define some auxiliary modules of interest. By the proof of [67, 9.4.1], both $\Delta:A\to A\otimes A$ and $\operatorname{id}\otimes S:A\otimes A\to A^{e}$ are algebra morphisms; write $\delta$ for their composition. We let $\mathcal{R}$ be the $A^{e}$-$A$-bimodule given by taking the diagonal $A^{e}$-bimodule and restricting the right action along $\delta$. Concretely, using Sweedler notation, the action is given by the formula $(a\otimes b).(x\otimes y).c=axc_{1}\otimes S(c_{2})yb$. Clearly $\mathcal{R}$ is free as a left $A^{e}$-module. Similarly, we write $\delta^{\prime}$ for the composition $(S\otimes\operatorname{id})\Delta$, and $\mathcal{L}$ for the analogous $A$-$A^{e}$-bimodule. We will sometimes wish to view $\mathcal{R}$ and $\mathcal{L}$ as $A$-bimodules, in which case we forget part of the $A^{e}$-action.
The following is our key intermediate lemma (cf. [35, 4.12] for a similar statement):
@@
We will only prove the desired statements for $\mathcal{R}$; the analogous statements for $\mathcal{L}$ then follow by symmetry. Claim (1) follows from the proof of [67, 9.2.9], which requires that $S$ be bijective (concretely, $\mathcal{R}$ is the free right $A$-module on the vector space $A$). Claim (2) follows from Claim (1) together with the fact that $\mathcal{R}\otimes_{A}k\cong A$, which is proved in [67, 9.4.2] (and does not require any condition on the antipode). For claim (3), since $\mathcal{R}$ is free of rank one as a left $A^{e}$-module, we see that the $A$-$A^{e}$-bimodule ${\mathrm{\mathbb{R}}}\mathrm{Hom}_{{A^{e}}}(\mathcal{R},A^{e})$ is simply given by $A^{e}$ where the right action is the regular $A^{e}$-action and the left action is given through the antipode $S$. But this is precisely $\mathcal{L}$, as desired. ∎
@@
$\simeq{\mathrm{\mathbb{R}}}\mathrm{Hom}_{A^{e}}(\mathcal{R}\otimes^{\mathbb{L}}_{A}k,A^{e})$ by 9.11(2)
@@
$\simeq{\mathrm{\mathbb{R}}}\mathrm{Hom}_{A^{\circ}}(k,\mathcal{L})$ by 9.11(3)
@@
$A^{*}$ $\simeq(\mathcal{R}\otimes^{\mathbb{L}}_{A}k)^{*}$ by 9.11(2)
@@
Since $\mathfrak{g}$ is finite dimensional, $U(\mathfrak{g})$ has finite global dimension (equal to the dimension of $\mathfrak{g}$). The equivalence $(1)\iff(2)$ now follows from [36, 2.3] (ultimately, this result goes back to Hazewinkel [31]). Since $U(\mathfrak{g})$ has finite global dimension, it is regular, and hence $(2)\iff(3)$ follows from 9.8. Since we have $CE(\mathfrak{g})\simeq B(U(\mathfrak{g}))$, the equivalences $(2)\iff(5)$ and $(3)\iff(4)$ follow from 8.3. ∎
@@
A more structured version of this example appears as [35, Theorem 1.3]; since $U(\mathfrak{g})$ is regular one can lift the Poincaré duality class to a class in cyclic homology.
@@ 9.4. Poincaré duality spaces
The material in this section is inspired by [7, §5.1], which was itself inspired by [48]. Recall that a simplicial set $K$ is grouplike if its homotopy category $hK$ is a groupoid.
@@
Let $C_{\bullet}(K)$ be the coalgebra of simplicial chains on $K$; when $K$ is a topological space, $C_{\bullet}(K)$ is the coalgebra of singular chains on $K$. We will also write $C^{\bullet}(K)$ for the linear dual of $C_{\bullet}(K)$, regarded as a plain dg algebra with no pseudocompact structure. Let $G(K)$ denote the loop group of $K$; when $K$ is a path connected topological space, $G(K)$ is homotopy equivalent to the based loop space $\Omega K$ and we will interchange these notions freely. Since $G(K)$ is a simplicial group, $C_{\bullet}(G(K))$ is an algebra under composition of loops. Note that when $K$ is connected, $C_{\bullet}(G(K))$ is connective.
@@
Say that a topological space is finitely dominated if it is a retract (in the homotopy category) of a finite CW complex. A finitely dominated space is homotopy equivalent to a CW complex [30, A.11].
@@
A necessary condition for a CW complex $X$ to be finitely dominated is that its total integral homology $H_{*}(X,\mathbb{Z})$ is a finitely generated abelian group. If $X$ is simply connected then this condition is sufficient [30, 4C.1].
@@
By 8.3 combined with 9.19, $X$ is a Frobenius space if and only if $C_{\bullet}\Omega X$ is a Gorenstein algebra. Our main application is the following theorem, which shows that a sufficiently finite space is a Poincaré duality space if and only if it is a Frobenius space:
@@
    $C_{\bullet}\Omega(X)$ is a $d$-Gorenstein algebra.
@@
    $C_{\bullet}\Omega(X)$ is a $d$-Calabi–Yau algebra.
@@
The equivalence of (2) and (3) is definitional, and we have already observed that (3) is equivalent to (5). The equivalence of (1) and (5) follows from 9.3, since by definition $X$ is a Poincaré duality space if and only if $C_{\bullet}\Omega X$ has Gorenstein duality. Note that $C_{\bullet}\Omega X$ is a regular involutive dg Hopf algebra, with antipode $S$ given by taking the inverse of a loop. So the equivalence of (5) and (6) follows from 9.8. The equivalence of (6) with (4) follows from 8.3 combined with 9.19. ∎
@@
Although it is not possible to formulate our notion of Poincaré duality space for non-finitely dominated topological spaces, we may still study the other conditions of 9.30. Indeed, if $X$ is any path connected topological space then the conditions $(2)$, $(3)$ and $(5)$ are all equivalent. Moreover, $(4)$ is equivalent to $(6)$.
@@
To see this, follow the proof of 9.30 and use 9.8(2) instead of 9.8(1). We suggest that any of the above equivalent conditions are a replacement for the concept of Poincaré duality space for spaces that are not finitely dominated but do have finite loop homology. In this sense, spaces with finite loop homology are ‘Koszul dual’ to finitely dominated spaces.

### v2

@@ 5.3. Properness for coalgebras
Let $M\in\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)$ be an acyclic comodule. We wish to show that $M$ is coacyclic. Following [58, 5.5], we see that the full subcategory of $D^{\mathrm{co}}(C)$ consisting of the acyclic $C$-comodules is precisely the left orthogonal to $C$. But this is also the left orthogonal to $\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)$. So $M$ is simultaneously an object of both $\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)$ and its left orthogonal, and hence is zero in $D^{\mathrm{co}}(C)$, as required. ∎
@@
There is a natural localisation functor $D^{\mathrm{co}}(C)\to D({\mathbf{Comod}}\text{-}C)$ whose kernel consists of the acyclic comodules. Under the assumption of Vopěnka’s principle, there is a natural equivalence $\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)\simeq D({\mathbf{Comod}}\text{-}C)$ so that $\mathrm{Coloc}_{D^{\mathrm{co}}(C)}(C)$ is actually the largest subcategory of $D^{\mathrm{co}}(C)$ for which the natural functor to $D(k)$ reflects isomorphisms. See [58, 5.5, Note added three years later] for further discussion.
@@ 6.2. Calabi–Yau algebras
The original definition of Ginzburg CY in [23] requires in addition that the isomorphism $f:A\to A^{\vee}[n]$ in the derived category of $A$-bimodules satisfies the extra condition that the composition $A\to A^{\vee\vee}\xrightarrow{f^{\vee}[n]}A^{\vee}[n]$ is equal to $f$. However, it is a theorem of Van den Bergh that in the smooth setting, this extra condition is superfluous [68, C.1]. In the nonsmooth setting, this extra condition does not seem useful.
@@ 6.3. Calabi–Yau conditions on categories
If $A$ is Ginzburg $n$-CY then the pair $(D(A),{{\mathbf{pvd}}}\kern 1.0pt(A))$ has a $n$-CY structure [43, 3.4]. If $A$ is smooth, proper, and finitely presented, then $[n]$ is a Serre functor on ${{\mathbf{pvd}}}\kern 1.0pt(A)$ if and only if $A$ is Ginzburg $n$CY [23, 3.2.4].
@@
We follow the proof of [43, 4.1]. There are natural maps
@@ 7.2. Discrete Gorenstein algebras
Let $R$ be a commutative noetherian $k$-algebra of finite Krull dimension $d$. Recall from [39, 40] that $R$ is said to be singular CY if it is Gorenstein and equicodimensional [39, 3.10]. An algebra essentially of finite type is noetherian and of finite Krull dimension, and an irreducible algebra is equicodimensional. Hence, if $R$ is irreducible and essentially of finite type, then it is singular CY if and only if it is Gorenstein, and 7.8 provides implications
@@ 7.3. Frobenius algebras
Recall that a discrete finite dimensional graded algebra $A$ is $d$-Frobenius if there is an isomorphism $A^{*}\cong A[d]$ of $A$-modules (equivalently, $A^{\circ}$-modules). Such an algebra is equipped with a Nakayama automorphism $\nu$, and one has an $A$-bimodule isomorphism $A^{*}\cong A_{\nu}[d]$, where $A_{\nu}$ denotes $A$ with the usual bimodule structure on the left, and twisted by $\nu$ on the right. One then obtains a quasi-isomorphism $\operatorname{HH}_{\bullet}(A,A_{\nu})^{*}\simeq\operatorname{HH}^{\bullet}(A)[d]$. A symmetric Frobenius algebra is one whose Nakayama automorphism is the identity, so if $A$ is symmetric one has the simpler description $\operatorname{HH}_{\bullet}(A)^{*}\simeq\operatorname{HH}^{\bullet}(A)[d]$. For a comprehensive reference on the classical theory of Frobenius algebras, see [49, §16].
@@ 8.1. The main theorem
For all statements, the proof consists of matching up appropriate modules, co/contramodules, and functors, across the relevant version of Koszul duality. We begin with (1). By 3.9, across bimodule Koszul duality the diagonal bimodule $A$ corresponds to the diagonal bicomodule $C$. Hence by 3.15, across bimodule Koszul duality the inverse dualising complex $A^{\vee}$ corresponds to $D(C)$. So one is invertible if and only if the other is, which proves (1). The proof of (2) is identical.
@@
For (4), 3.1(6) shows that across one-sided Koszul duality the left $C$-contramodule $k$ corresponds to the left $A$-module $A^{*}$. Hence 3.18 shows that the left $C$-contramodule ${\mathrm{\mathbb{R}}}\mathrm{Hom}^{C}(k,C^{*})$ corresponds to the left $A$-module $A^{**}$. Hence if $A$ is $d$-Frobenius, we have $A^{*}\simeq A^{**}[d]$ as left $A$-modules and hence $C$ is $d$-Gorenstein. In the other direction, if $C$ is $d$-Gorenstein then the left $A$-module $M\coloneqq A^{*}$ is isomorphic to $M^{*}[d]$, which implies that $M\in{{\mathbf{ref}}}\kern 1.0pt{A}$ and hence $A\in{{\mathbf{ref}}}\kern 1.0pt{A}$. Hence $A^{**}\simeq A$ and so A is $d$-Frobenius.
For (5), 3.10 shows that across two-sided Koszul duality the bimodule $A^{*}$ corresponds to the bicontramodule $C^{*}$. Hence 3.20 implies that the $A$-bimodule $A^{**}$ corresponds to the $C$-bicontramodule $C^{\vee}$. If $A$ is twisted symmetric, we see that there is an invertible bimodule $X$ such that $A^{*}\simeq{}_{X}(A^{**})$. Translating this across Koszul duality, we see that there is an invertible bicomodule $Y$ such that $C^{*}\simeq{}_{Y}(C^{\vee})$; in other words $C$ is twisted CY. Conversely if $C$ is twisted CY then with we obtain an invertible $A$-bimodule $X$ and an $A$-bilinear quasi-isomorphism $A^{*}\simeq{}_{X}(A^{**})$, and so putting $M\coloneqq A^{*}$ we have $A$-bilinear quasi-isomorphisms
@@
and hence $R\otimes S$ is $(r+s)$-Gorenstein. For (2), suppose that $C$ is regular $r$-Frobenius and $D$ is regular $s$-Frobenius. By [34, 5.1] we have a natural algebra quasi-isomorphism $\Omega(C\otimes D)\simeq\Omega(C)\otimes\Omega(D)$. Putting $R\coloneqq\Omega C$ and $S\coloneqq\Omega D$ we see by 8.3 that $R$ is $r$-Gorenstein and $S$ is $s$-Gorenstein. Moreover $R$ is proper since $C$ is regular. Hence by (1), $R\otimes S$ is $(r+s)$-Gorenstein, and hence $C\otimes D$ is $(r+s)$-Frobenius by 8.3 again. ∎
@@ 8.2. Proper Frobenius coalgebras
Since $C$ is proper, it is weakly equivalent to a finite dimensional comodule over itself, and hence the comodule $C^{*}$ is the usual linear dual of $C$. By dualising, it follows that $A$ is $(-d)$-Frobenius if and only if $C$ is quasi-isomorphic to $C^{*}[d]$ as $C$-comodules. Since weak equivalences are quasi-isomorphisms, $(1)$ is now clear. The second claim follows similarly from 8.12. The third claims are proved identically; for (4) only needs to observe in addition that if $C$ is simply connected then so is $C^{e}$. ∎
@@
The simply connectedness assumptions of the above proposition cannot be dropped. Counterexamples can be found in chain coalgebras of topological spaces; see 9.43 for the details.
@@ 9.3. Hopf algebras
If $A$ is a dg Hopf algebra, and $A^{\prime}$ is a dg algebra with an algebra quasi-isomorphism $A^{\prime}\simeq A$, then $A^{\prime}$ is Calabi–Yau or Gorenstein precisely when $A$ is. In particular we may apply the above theorem to $A^{\prime}$ verbatim. We will make use of this observation later.
Remark 9.10.
@@
To begin the proof of 9.8, we need to define some auxiliary modules of interest. By the proof of [69, 9.4.1], both $\Delta:A\to A\otimes A$ and $\operatorname{id}\otimes S:A\otimes A\to A^{e}$ are algebra morphisms; write $\delta$ for their composition. We let $\mathcal{R}$ be the $A^{e}$-$A$-bimodule given by taking the diagonal $A^{e}$-bimodule and restricting the right action along $\delta$. Concretely, using Sweedler notation, the action is given by the formula $(a\otimes b).(x\otimes y).c=axc_{1}\otimes S(c_{2})yb$. Clearly $\mathcal{R}$ is free as a left $A^{e}$-module. Similarly, we write $\delta^{\prime}$ for the composition $(S\otimes\operatorname{id})\Delta$, and $\mathcal{L}$ for the analogous $A$-$A^{e}$-bimodule. We will sometimes wish to view $\mathcal{R}$ and $\mathcal{L}$ as $A$-bimodules, in which case we forget part of the $A^{e}$-action.
The following is our key intermediate lemma (cf. [37, 4.12] for a similar statement):
@@
We will only prove the desired statements for $\mathcal{R}$; the analogous statements for $\mathcal{L}$ then follow by symmetry. Claim (1) follows from the proof of [69, 9.2.9], which requires that $S$ be bijective (concretely, $\mathcal{R}$ is the free right $A$-module on the vector space $A$). Claim (2) follows from Claim (1) together with the fact that $\mathcal{R}\otimes_{A}k\cong A$, which is proved in [69, 9.4.2] (and does not require any condition on the antipode). For claim (3), since $\mathcal{R}$ is free of rank one as a left $A^{e}$-module, we see that the $A$-$A^{e}$-bimodule ${\mathrm{\mathbb{R}}}\mathrm{Hom}_{{A^{e}}}(\mathcal{R},A^{e})$ is simply given by $A^{e}$ where the right action is the regular $A^{e}$-action and the left action is given through the antipode $S$. But this is precisely $\mathcal{L}$, as desired. ∎
@@
$\simeq{\mathrm{\mathbb{R}}}\mathrm{Hom}_{A^{e}}(\mathcal{R}\otimes^{\mathbb{L}}_{A}k,A^{e})$ by 9.12(2)
@@
$\simeq{\mathrm{\mathbb{R}}}\mathrm{Hom}_{A^{\circ}}(k,\mathcal{L})$ by 9.12(3)
@@
$A^{*}$ $\simeq(\mathcal{R}\otimes^{\mathbb{L}}_{A}k)^{*}$ by 9.12(2)
@@
Since $\mathfrak{g}$ is finite dimensional, $U(\mathfrak{g})$ has finite global dimension (equal to the dimension of $\mathfrak{g}$). The equivalence $(1)\iff(2)$ now follows from [38, 2.3] (ultimately, this result goes back to Hazewinkel [32]). Since $U(\mathfrak{g})$ has finite global dimension, it is regular, and hence $(2)\iff(3)$ follows from 9.8. Since we have $CE(\mathfrak{g})\simeq B(U(\mathfrak{g}))$, the equivalences $(2)\iff(5)$ and $(3)\iff(4)$ follow from 8.3. ∎
@@
A more structured version of this example appears as [37, Theorem 1.3]; since $U(\mathfrak{g})$ is regular one can lift the Poincaré duality class to a class in cyclic homology.
@@ 9.4. Poincaré duality spaces
The material in this section is inspired by [7, §5.1], which was itself inspired by [50]. Similar results appear in [37, §4.2]. Recall that a simplicial set $K$ is grouplike if its homotopy category $hK$ is a groupoid.
@@
Let $C_{\bullet}(K)$ be the coalgebra of simplicial chains on $K$; when $K$ is a topological space, $C_{\bullet}(K)$ is the coalgebra of singular chains on $K$. We will also write $C^{\bullet}(K)$ for the linear dual of $C_{\bullet}(K)$, regarded as a plain dg algebra with no pseudocompact structure. Let $G(K)$ denote the Kan loop group of $K$; when $K$ is (the singular simplicial set of) a path connected topological space, $GK$ is homotopy equivalent to the based loop space $\Omega K$ and we will interchange these notions freely. Since $G(K)$ is a simplicial group, $C_{\bullet}(G(K))$ is an algebra under composition of loops. In fact, $C_{\bullet}(G(K))$ is an involutive dg Hopf algebra; the coalgebra structure is that of simplicial chains and the antipode is given by reversing loops. Note that when $K$ is connected, $C_{\bullet}(G(K))$ is connective.
Remark 9.20.
If $X$ is a topological space and one uses Moore loops, then $C_{\bullet}(\Omega X)$ becomes a dg algebra, and there is a quasi-isomorphism $C_{\bullet}(\Omega X)\simeq C_{\bullet}(G\mathrm{Sing}X)$ of dg algebras. In the sequel we will tacitly assume that $C_{\bullet}(\Omega X)$ is a dg algebra in this manner.
@@
Say that a topological space is finitely dominated if it is a retract (in the homotopy category) of a finite CW complex. A finitely dominated space is homotopy equivalent to a CW complex [31, A.11].
@@
A necessary condition for a CW complex $X$ to be finitely dominated is that its total integral homology $H_{*}(X,\mathbb{Z})$ is a finitely generated abelian group. If $X$ is simply connected then this condition is sufficient [31, 4C.1].
@@
By 8.3 combined with 9.21, $X$ is a Frobenius space if and only if $C_{\bullet}\Omega X$ is a Gorenstein algebra. Our main application is the following theorem, which shows that a sufficiently finite space is a Poincaré duality space if and only if it is a Frobenius space:
@@
    $C_{\bullet}\Omega X$ is a $d$-Gorenstein algebra.
@@
    $C_{\bullet}\Omega X$ is a $d$-Calabi–Yau algebra.
@@
The equivalence of (2) and (3) is definitional, and we have already observed that (3) is equivalent to (5). The equivalence of (1) and (5) follows from 9.3, since by definition $X$ is a Poincaré duality space if and only if $C_{\bullet}\Omega X$ has Gorenstein duality. By 9.9 and 9.20 we may assume that $C_{\bullet}\Omega X$ is an involutive dg Hopf algebra, which is necessarily regular by 9.29. So the equivalence of (5) and (6) follows from 9.8. The equivalence of (6) with (4) follows from 8.3 combined with 9.21. ∎
@@
Although it is not possible to formulate our notion of Poincaré duality space for non-finitely dominated topological spaces, we may still study the other conditions of 9.32. Indeed, if $X$ is any path connected topological space then the conditions $(2)$, $(3)$ and $(5)$ are all equivalent. Moreover, $(4)$ is equivalent to $(6)$.
@@
To see this, follow the proof of 9.32 and use 9.8(2) instead of 9.8(1). We suggest that any of the above equivalent conditions are a replacement for the concept of Poincaré duality space for spaces that are not finitely dominated but do have finite loop homology. In this sense, spaces with finite loop homology are ‘Koszul dual’ to finitely dominated spaces.